A drug is 63% bioavailable by the intramuscular route and 42% bioavailable by the oral route. How many 150 mg tablets would need to be administered b.i.d. to receive the equivalent total intramuscular dose that delivers 8 mg/hr
Answer #1

The drug's bioavailability by intramuscular route=63%

But bioavailability by oral route = 42%

So IM dose of 8 ml/hr = 8 * 24= 192 mg

So 192 mg is available by IM route and must be available by oral route

So 42% of the oral dose administered = 192 mg

Oral dose = 192 * 100 / 42 =457.15 mg

Bid means twice daily.

So per dose = 457.15/2 = 228.6 mg

So one and a half tablets of 150 mg should be administered BID.